3. Divergence Find the divergence of: a) Ē(x, y, z)=(-2y x 0] b) F(x,y,z)= (y2 – 2x 5x’y x+37] c) v =[3y–2yx xy? -62?x]
Answer #1

The divergence of

\vec{F}(x,y,z)=\left ( P(x,y,z),Q(x,y,z),R(x,y,z) \right )

div.\vec{F}(x,y,z)=\bigtriangledown .\vec{F}

div.\vec{F}(x,y,z)=\frac{\partial P}{\partial x}+\frac{\partial Q}{\partial y}+\frac{\partial R}{\partial z}

a)

P = -2y

Q = x

R = 0

div.\vec{F}(x,y,z)=\frac{\partial (-2y)}{\partial x}+\frac{\partial (x)}{\partial y}+\frac{\partial (0)}{\partial z}

div.F(x, y, z) = 0+0+0=0

div.\vec{F}(x,y,z)=0

b)

P = y2-2x

Q = 5x2y

R = x+3z

div.\vec{F}(x,y,z)=\frac{\partial (y^{2}-2x)}{\partial x}+\frac{\partial (5x^{2}y)}{\partial y}+\frac{\partial (x+3z)}{\partial z}

div.\vec{F}(x,y,z)= -2 + 5x^{2}+3 = 5x^{2}+1

div.\vec{F}(x,y,z)= 5x^{2}+1

c)

P = 3y-2yx

Q = xy2

R = -6z2x

div.\vec{F}(x,y,z)=\frac{\partial (3y-2yx)}{\partial x}+\frac{\partial (xy^{2})}{\partial y}+\frac{\partial (-6z^{2}x)}{\partial z}

div.\vec{F}(x,y,z)= -2y+2xy-12xz

Note : If you have any doubt feel free to comment. I will reply ASAP. If you find my answer useful please give me a thumbs up.It will be a huge help for beginners like me.